OpenStudy (anonymous):

use the given values to evaluate all six trigonometric functinons A. cot theta = -3, sin theta <0 B. cos x= -3/5, tan x=4/3

OpenStudy (anonymous):

the formula you need here is sin^2+cos^2=1

OpenStudy (anonymous):

how would i plug the given numbers in?

OpenStudy (anonymous):

I see on the first problem that my method won't work, so let's use another one e.mccormick pm'd me this: On heythereee's problem.... why not use the facts that sin=y/r, cos=x/r, and tan=y/x to find the values of x, y, and r. Then fill out the other four trig ratios.

OpenStudy (anonymous):

hmm. i'll try that